# Unlearning Isn’t Deletion: Investigating Reversibility of Machine Unlearning in LLMs

### v1

@@ 6 Conclusion
We revisit machine unlearning for LLMs through a systematic study of reversibility. Token-level metrics alone can mislead: a model may appear collapsed yet remain fully recoverable.
To diagnose this gap, we introduce a feature-space toolkit—PCA Similarity, PCA Shift, CKA, and FIM—that cleanly separates reversible from irreversible catastrophic forgetting. Our empirical and theoretical results show that true forgetting arises only when many layers undergo coordinated, large-magnitude perturbations; by contrast, minor updates in high-sensitivity regions (e.g., output logits) can slash accuracy or inflate perplexity while leaving internal representations intact.
These findings call for evaluation protocols that go beyond token-level scores and for algorithms that actively control representational drift. We further observe that unlearning, followed by proper relearning, can refine representations and even boost downstream performance. Together, these insights chart a path toward safer, more interpretable unlearning in LLMs.
@@ A.1 Limitations
Our experiments target two LLMs and a handful of tasks and unlearning methods; although our diagnostic framework is model-agnostic and designed to scale, empirical validation on much larger models and production-scale pipelines remains to be done. The constrained relearning protocol and selected metrics provide clear insights into representational drift but are not exhaustive and do not offer formal privacy guarantees. Extending these analyses to diverse architectures, richer unlearning strategies, and rigorous theoretical bounds is an important direction for future work.
A.2 Broader impacts
This work provides new tools and theoretical insights for evaluating and guiding machine unlearning in LLMs, with the potential to improve privacy guarantees and regulatory compliance (e.g., GDPR “right to be forgotten”). By distinguishing reversible from irreversible forgetting, our diagnostics can help developers ensure that sensitive or copyrighted material is truly expunged rather than merely hidden. At the same time, adversaries could exploit reversible forgetting pathways to mask malicious or biased content and then restore it later, highlighting the need for robust defenses. The computational overhead of continual unlearning and repeated diagnostics may increase energy consumption, underscoring the importance of efficient implementations. Finally, by exposing the fragility of token-level metrics, our work advocates for more trustworthy evaluation standards that balance privacy, utility, and environmental considerations in real-world deployments.
@@ A.3 Detailed Analysis Tools
$\operatorname{PCA\text{-}Sim}(i)=\cos\bigl(\mathbf{c}_{i,1}^{\text{orig}},\mathbf{c}_{i,1}^{\text{upd}}\bigr{missing})=\frac{(\mathbf{c}_{i,1}^{\text{orig}})^{\top}\mathbf{c}_{i,1}^{\text{upd}}}{\lVert\mathbf{c}_{i,1}^{\text{orig}}\rVert\,\lVert\mathbf{c}_{i,1}^{\text{upd}}\rVert}\in[-1,1],$
where values near $1$ indicate stable directional alignment, and values near $-1$ suggest a near-orthogonal shift in dominant directions.
@@
$\operatorname{PCA\text{-}Shift}(i)=p_{1,\text{upd}}-p_{1,\text{orig}},\qquad\operatorname{Principle}(i)=p_{2,\text{upd}},$
where $\operatorname{PCA\text{-}Shift}$ quantifies displacement along PC1 and $\operatorname{Principle}$ captures orthogonal deviation along PC2. These metrics reflect how the representation center drifts within the top subspace.
@@
where values near $1$ indicate highly overlapping subspaces, and values near 00 signal near-orthogonality.
@@ A.4 Complete Experiment
Table 4 summarizes token-level performance under four unlearning configurations on Yi-6B, varying both learning rates and request counts. Most metrics—including forget/retain accuracy, perplexity, CSQA, and GSM8K—follow consistent trends: milder methods (e.g., NPO, NPO+KL) yield moderate degradation and support effective recovery via relearning, while aggressive methods (e.g., GA, RL) under high learning rates lead to severe, often irrecoverable performance drops. These results align closely with our theoretical distinction: reversible (catastrophic) forgetting corresponds to localized perturbations and structurally recoverable models, while irreversible (catastrophic) forgetting reflects widespread, irreversible damage.
MIA, however, behaves differently. Even when downstream performance collapses, MIA scores often remain high and recover rapidly. This suggests MIA is especially sensitive to shallow changes (e.g., output logits or normalization), rather than deeper representational shifts. As a result, MIA often correlates with reversible forgetting—capturing surface-level instability without revealing structural collapse. These observations reinforce that while token-level metrics offer partial signals, only internal diagnostics can reliably differentiate reversible from irreversible forgetting.
A.5 Single Unlearning
Figure 6 illustrates feature-level changes under single unlearning using PCA Similarity, PCA Shift, CKA, and Fisher Information. In (a), PCA Similarity remains consistently high across layers, with cosine scores near 1, indicating that dominant activation directions are well preserved. Slight dips in shallow and final layers are quickly restored after relearning, suggesting minimal and reversible geometric drift. Subfigure (b) confirms that PC1 shifts and orthogonal deviations are small, with relearned centers closely matching the original. In (c), CKA shows near-perfect alignment between the original, unlearned, and relearned representations, reinforcing the conclusion that subspace structure remains intact. Fisher Information spectra in (d–f) reveal only mild leftward shifts, indicating slight loss flattening and reduced parameter sensitivity, which are fully recovered after relearning. Overall, these results confirm that single unlearning causes only minor, reversible structural perturbations—highlighting the fragility of token-level evaluations in capturing irreversible forgetting.
(a) Simple (Single Unlearning)
(b) Simple (Single Unlearning)
(c) Simple (Single Unlearning)
(d) Simple (Single Unlearning)
(e) Simple (Single Unlearning)
(f) Simple (Single Unlearning)
Figure 6: Single unlearning analysis on Yi-6B for GA under a simple task. PCA Similarity, PCA Shift, CKA, and Fisher information (FIM) distributions are reported across key layers to assess representation and parameter changes before and after unlearning and relearning.
A.6 Detailed Analysis Results
@@
[…]
@@
[…]
@@
[…]
@@
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
@@
(a) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple Task ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple Task ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
[…]
(a) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple Task ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
@@
(a) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex Task ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex Task ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex Task ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
[…]
@@
(a) LR=$3\times 10^{-6}$, Layer 28
(b) LR=$5\times 10^{-6}$, Layer 28
(c) LR=$3\times 10^{-5}$, Layer 28
(d) LR=$3\times 10^{-6}$, Layer 22
(e) LR=$5\times 10^{-6}$, Layer 22
(f) LR=$3\times 10^{-5}$, Layer 22
(g) LR=$3\times 10^{-6}$, Layer 13
(h) LR=$5\times 10^{-6}$, Layer 13
(i) LR=$3\times 10^{-5}$, Layer 13
(j) LR=$3\times 10^{-6}$, Layer 4
(k) LR=$5\times 10^{-6}$, Layer 4
(l) LR=$3\times 10^{-5}$, Layer 4
(m) LR=$3\times 10^{-6}$, Layer 1
(n) LR=$5\times 10^{-6}$, Layer 1
(o) LR=$3\times 10^{-5}$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1
@@
(a) LR=$3\times 10^{-6}$, Layer 31
(b) LR=$5\times 10^{-6}$, Layer 31
(c) LR=$3\times 10^{-5}$, Layer 31
(d) LR=$3\times 10^{-6}$, Layer 28
(e) LR=$5\times 10^{-6}$, Layer 28
(f) LR=$3\times 10^{-5}$, Layer 28
(g) LR=$3\times 10^{-6}$, Layer 22
(h) LR=$5\times 10^{-6}$, Layer 22
(i) LR=$3\times 10^{-5}$, Layer 22
(j) LR=$3\times 10^{-6}$, Layer 13
(k) LR=$5\times 10^{-6}$, Layer 13
(l) LR=$3\times 10^{-5}$, Layer 13
(m) LR=$3\times 10^{-6}$, Layer 4
(n) LR=$5\times 10^{-6}$, Layer 4
(o) LR=$3\times 10^{-5}$, Layer 4
(p) LR=$3\times 10^{-6}$, Layer 1
(q) LR=$5\times 10^{-6}$, Layer 1
(r) LR=$3\times 10^{-5}$, Layer 1

### v2

@@ 6 Conclusion
This work demonstrates that class (task-level) evaluations of LLM unlearning are insufficient, as performance collapse often masks the reversibility of forgetting. Models may appear to have erased data while their internal representations remain intact and easily recoverable. Our representation-level toolkit reveals that genuine forgetting requires substantial, coordinated weight perturbations. Minor updates often create only a superficial, reversible effect. We find that achieving the ideal goal of irreversible, non-catastrophic forgetting remains an open challenge, exposing a fundamental limitation in current methods. Our findings call for a shift in evaluation, moving beyond surface-level metrics to protocols that measure true representational change. This is essential for developing unlearning algorithms that can provide meaningful and trustworthy guarantees of data removal.
@@ A.1 Limitations
Our experiments target two LLMs and a handful of tasks and unlearning methods; although our diagnostic framework is model-agnostic and designed to scale, empirical validation on much larger models and production-scale pipelines remains to be done. The constrained relearning protocol and selected metrics provide clear insights into representational drift but are not exhaustive and do not offer formal privacy guarantees. In this work, we primarily relied on four diagnostic tools—PCA similarity, PCA shift, CKA, and Fisher information—to capture different aspects of representational drift. Other feature-level methods, such as correlation-based approaches (e.g., SVCCA [30]), offer similar perspectives on subspace similarity. Incorporating a broader suite of analytic tools is an important direction for future work.
A.2 Related Work
Machine Unlearning.
Machine unlearning has emerged as a critical direction for addressing privacy, safety, and bias in large language models (LLMs) [40; 12; 29; 20; 23; 7; 32; 37; 44; 42; 38; 36; 2]. It is typically defined as either exact or approximate [2]. Exact unlearning requires the resulting model to be indistinguishable from one retrained from scratch on the retain set, fully eliminating any statistical trace of the forget set. Approximate unlearning relaxes this requirement to distributional or behavioral similarity, demanding only comparable outputs (e.g., perplexity or forget-set accuracy) between unlearned and retrained models [27; 32]. For modern LLMs, however, exact unlearning is computationally infeasible, as full retraining or partition-based schemes scale poorly [2]. Consequently, approximate methods dominate practice in LLMs.
Single-Shot Unlearning.
Most existing approaches are designed for single deletion events. Gradient-based strategies (e.g., GA) enforce forgetting directly but often incur significant utility loss [40]. Recent advances such as WAGLE augment these methods with weight attribution (e.g., GA+GD+WAGLE), selectively updating the most influential parameters to enhance forgetting efficacy while mitigating utility degradation [13]. Prompt-based steering avoids parameter updates, reducing cost, but typically achieves only superficial forgetting with vulnerability to reactivation [23]. Model-editing methods, such as AlphaEdit [21], are lightweight and potentially robust, yet their behavior under sequential or heterogeneous requests remains underexplored.
Continual Unlearning.
When unlearning requests arrive sequentially, naive extensions of single-shot methods tend to compound damage, leading to catastrophic forgetting and unstable dynamics [1; 32]. Each request operates on an already modified model, magnifying utility loss. Recent work has attempted to mitigate this through orthogonal updates (e.g., LoRA-based unlearning [11]) and OOD detectors. ALKN [36] advances this line by providing a principled framework for continual unlearning, introducing parameter-level interventions and adaptive modules to counteract accumulative decline.
Evaluations.
Evaluating unlearning efficacy remains an open challenge. Existing studies rely on three main classes of metrics. First, classic (task-level) metrics such as accuracy, perplexity [40; 20] are widely used but can be misleading, since performance degradation does not guarantee removal of knowledge. Second, memorization probes [17] assess verbatim recall, offering finer granularity but failing to capture semantic or paraphrased knowledge. Third, robustness-based evaluations examine vulnerabilities to jailbreaking [46], relearning attacks [24], and even quantization attacks [44]. For quantization attacks, low-bit compression restores forget-set behavior without direct access to the forget. While existing approaches expose weaknesses in reversibility, they often conflate forgetting with class task-level degradation and lack structural insight. Our representation-level toolkit closes this gap by jointly diagnosing reversibility and catastrophicity, yielding a more faithful understanding of what is truly forgotten. We apply this toolkit to both single unlearning and continual unlearning, the latter of which has not been systematically investigated despite being a more realistic scenario where deletion requests arrive sequentially over a model’s lifecycle.
@@ A.3 Detailed Analysis Tools
$\operatorname{PCA\text{-}Sim}(i)=\cos\bigl(\mathbf{c}_{i,1}^{\text{orig}},\mathbf{c}_{i,1}^{\text{upd}}\bigr.)=\frac{(\mathbf{c}_{i,1}^{\text{orig}})^{\top}\mathbf{c}_{i,1}^{\text{upd}}}{\lVert\mathbf{c}_{i,1}^{\text{orig}}\rVert\,\lVert\mathbf{c}_{i,1}^{\text{upd}}\rVert}\in[-1,1],$
where values near $1$ indicate stable directional alignment, and values near 1-1 suggest a near-orthogonal shift in dominant directions.
@@
${PC1\ \Delta}(i)={\mathbf{c}_{i,1}}^{\text{upd}}-{\mathbf{c}_{i,1}}^{\text{orig}},\qquad{PC2}(i)={\mathbf{c}_{i,2}}^{\text{upd}},\qquad p_{i,12}=(PC1\ \Delta(i),PC2(i))$
where $p_{i,12}$} quantifies displacement along PC1 and captures orthogonal deviation along PC2. These metrics reflect how the representation center drifts within the top subspace.
@@
where values near $1$ indicate highly overlapping subspaces, and values near 0 signal near-orthogonality.
@@
Table 7: Qwen-2.5-7B: MIA / MATH / GSM8K Accuracy (%) for complex task under four settings. Bold numbers indicate improvements over the Original baseline in MATH or GSM8K.
[…]
MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K MIA MATH GSM8K
Original —– 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1 99.3 9.0 80.1
Unlearn GA 5.9 0.0 0.0 0.9 0.0 0.0 3.8 0.0 0.0 5.5 0.0 0.0
NPO 95.9 0.0 0.2 97.4 21.5 74.1 67.4 24.1 71.8 94.7 0.0 0.4
RLabel 35.5 0.0 0.0 69.6 0.0 1.5 11.2 0.0 0.0 2.9 0.0 0.0
Relearn GA 97.6 0.0 1.1 99.3 5.1 83.2 99.4 9.3 77.8 99.2 0.0 0.0
NPO 95.8 0.0 0.0 99.4 4.7 82.6 99.4 16.5 75.7 99.2 0.0 0.0
RLabel 99.5 0.0 0.0 99.3 5.3 83.3 99.3 10.0 77.2 99.6 0.0 0.0
Table 8: Yi-6B (GA): Mean PCA distance under different learning rates. The left block uses China Taiwan for relearning only, while the right block uses TOFU for both unlearning and relearning.
Relearning with China Taiwan Unlearning + Relearning with TOFU
Learning Rate Phase Mean PCA distance (forget set) Phase Mean PCA distance (forget set)
$3\times 10^{-6}$ Unlearn 17.12 Unlearn 0.51
$3\times 10^{-6}$ Relearn 4.98 Relearn 0.27
$5\times 10^{-6}$ Unlearn 20.27 Unlearn 2.41
$5\times 10^{-6}$ Relearn 10.77 Relearn 1.08
$3\times 10^{-5}$ Unlearn 193.13 Unlearn 11.96
$3\times 10^{-5}$ Relearn 167.32 Relearn 11.02
A.4 Mean PCA distance under Different Dataset
To examine the role of distributional alignment, we evaluate unlearning and relearning under two dataset settings. First, we use the TOFU benchmark [27], where both unlearning and relearning occur within the same distribution. Second, treating different languages as out-of-distribution (OOD), we include a Traditional-Chinese corpus for relearning. This setup enables us to probe whether cross-lingual signals can drive effective recovery, and how their efficacy compares with in-distribution.
Table 8 confirms that cross-lingual relearning improves the model but achieves less complete restoration than English data: mean PCA distance and related summary metrics move closer to baseline values, yet remain substantially higher. Greater linguistic or domain dissimilarity therefore reduces the efficacy of recovery, though partial restoration is still attainable.
For the TOFU dataset, the overall pattern holds: learning rate and the number of unlearning requests ($N$) effectively regulate feature drift and reversibility. However, the representational shifts induced by TOFU are milder than those observed in our simple and complex tasks. We attribute this to the smaller and less diverse nature of TOFU’s corpus; many entries are short and contain only author metadata, making its impact on the model’s feature space comparatively limited.
A.5 Detailed Analysis Results
A.5.1 Principal Component Analysis: Similarity and Shift
Across the same hyper-parameter grid, Figure 7 (PCA–Similarity) and Figure 11 (PCA–Shift) provide complementary views of representational drift. For GA, higher learning rates drive unlearned states (orange) far from the original (blue), while relearning (green) fails to return, producing long rays of irreversible drift. GA+GD narrows the spread but still collapses at $3{\times}10^{-5}$.
On Qwen-2.5-7B, GA shifts span thousands of PC1 units and drive PC2 to extreme negatives (Figure 13c,f,i), consistent with the multi-layer perturbations predicted in Section 4. In complex tasks such as mathematical reasoning, even small perturbations in hidden states can lead to substantial performance differences. This is reflected in our PCA–Similarity analysis, where seemingly minor changes in hidden state geometry correspond to meaningful behavioral variations. Besides, PCA-Similarity captures global alignment, whereas PCA–Shift highlights fine-grained translational drift. This distinction also explains why Figure 9h,i show only moderate misalignment under similarity but reveal pronounced displacements under shift (cf. Figure 13). Using both metrics thus provides a more complete characterization of reversibility. Overall, these results confirm that GA, with or without GD or KL, induces large and often irreversible displacements, whereas NPO variants, and to a lesser extent RLabel, constrain less shifts, consistent with our utility findings.
A.5.2 Centered Kernel Alignment Analysis
Figures 1517 report layer-wise linear CKA between the original model and its unlearned or relearned counterparts. Across both Yi-6B and Qwen-2.5-7B, GA stands out: as the learning rate or $N$ increases, its CKA curve drops close to zero in most layers and fails to recover, revealing a deep subspace fracture consistent with the irreversible PCA trends. GA+GD and GA+KL mitigate this decline to some extent but do not restore full alignment after relearning.
Task complexity does not alter the ordering but amplifies the differences. On the math-heavy Qwen benchmark, GA’s tail layers fall almost to zero at high learning rates, whereas NPO maintains significantly higher alignment. Taken together with the PCA-Shift results, these findings show that GA-style objectives consistently break subspace alignment, NPO variants preserve much greater stability, and RLabel induces moderate but partly recoverable distortions.
A.5.3 Fisher Information Analysis
Figures 1933 plot the empirical Fisher spectra layer by layer. Across both Yi-6B (simple) and Qwen-2.5-7B (complex), GA and its variants exhibit a pronounced leftward shift of the diagonal histogram as LR or $N$ increase. The peaks move several orders of magnitude in middle and deep layers, reflecting a flattened loss surface and diminished parameter salience. Crucially, these shifts persist after relearning, marking the onset of irreversible forgetting.
NPO, NPO+KL, and RL produce smaller leftward displacements under moderate LR or $N$, and their Fisher spectra recenter after relearning, indicating primarily reversible drift. Under extreme settings (e.g., ${\rm LR}=3\times 10^{-5}$ or $N=100$), these methods also show persistent displacement in some layers, suggesting milder but still irreversible forgetting.
Figures 14, 10, 18, and 34 examine relearning dynamics when the fine-tuning data and input query are drawn from the forget set, the retain set, or an unrelated data: i) across all sources, the overall trends are similar: alignment can be partially restored, but recovery is consistently weaker with unrelated data, underscoring that effective relearning depends on both the size and the relevance of the training set; ii) the observed behavior also varies with the choice of input queries. In the case of reversible catastrophic forgetting, all forget set, retain set, and unrelated data undergo the similar feature drifts.
(a) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 5: Layer-wise PCA Similarity and Shift for GA on Yi-6B (simple task). vary $N\in\{6,50,100\}$ at LR $3\times 10^{-5}$. Sustained low similarity or large shifts signal severe, irreversible catastrophic forgetting, whereas partial similarity or small shifts indicate mild, reversible catastrophic forgetting. Input queries are drawn from the forget set.
(a) ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) $N=6$, Layer 31
(e) $N=50$, Layer 31
(f) $N=100$, Layer 31
Figure 6: CKA and FIM for GA on Yi-6B, simple task. Vary LR $3\times 10^{-5}$ with $N\in\{6,50,100\}$. High CKA ( 1) and concentrated FIM spectra indicates reversible catastrophic forgetting, while persistently low CKA and large-shifted, flattened spectra denote severe representational drift and irreversible catastrophic forgetting. Input queries are drawn from the forget set.
@@ A.5.3 Fisher Information Analysis
Figure 7: PCA Similarity Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
@@
Figure 9: PCA Similarity Across Layers. Each row shows results under different unlearning methods: GA (a-c) NPO (d–f), Rlable (g–j). All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
@@
Figure 11: PCA Shift Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
@@
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
Figure 13: PCA Shift Across Layers. Each row shows results under different unlearning methods: GA (a-c) NPO (d–f), Rlable (g–j). All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.
@@
(a) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(b) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(e) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(h) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(k) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-6},\,N=100$)
(n) Simple ($\mathrm{LR}=5\times 10^{-6},\,N=100$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
Figure 15: CKA Across Layers. Each row shows results under different unlearning methods: GA+GD (a–c), GA+KL (d–f), NPO (g–i), NPO+KL (j–l), and Rlable (m–o). All plots are for the simple task on Yi-6B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=100$.
(a) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(b) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(c) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(d) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(e) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(f) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(g) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(h) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(i) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(j) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(k) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(l) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
(m) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(n) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=50$)
(o) Simple ($\mathrm{LR}=3\times 10^{-5},\,N=100$)
@@
(a) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(b) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(c) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(d) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(e) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(f) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
(g) Complex ($\mathrm{LR}=3\times 10^{-6},\,N=6$)
(h) Complex ($\mathrm{LR}=5\times 10^{-6},\,N=6$)
(i) Complex ($\mathrm{LR}=3\times 10^{-5},\,N=6$)
Figure 17: CKA Across Layers. Each row shows results under different unlearning methods: GA (a-c) NPO (d–f), Rlable (g–j). All plots are for the complex task on Qwen2.5-7B, using three learning rates $\{3\times 10^{-6},\,5\times 10^{-6},\,3\times 10^{-5}\}$ and fixed $N=6$.